### v1

@@ 5.2 Estimate of p
Figure 6: Distribution of the slope $p$ of the non-thermal population of accelerated particles for the sample of 22 GRBs analyzed in this work. This distribution has been derived from the $\beta$ distribution shown in Fig. 4, by using the relation $p=-2-2\beta$ or, equivalently, $p=2\beta_{e}$ (i.e. assuming synchrotron emission in the fast cooling regime). The median value of $p$ along with its error bar on $p$ values are represented with the black dot and solid line, respectively. For comparison, the distribution of $p$ derived for the Fermi/GBM Catalog bursts under the same assumption is also plotted (red dashed histogram).
There are two main considerations that we can draw from these results. The first one regards the width of the $p$ distribution. Fig. 6 shows that the distribution of $p$ is quite broad, with a $\sigma=1.12$. The majority of $p$ values are clustered between 2 and 3.5, but there is also a non-negligible tail extending towards steep values. This result may suggest that the acceleration mechanism does not produce relativistic particles distributed as a power–law with a unique slope, but rather with a quite wide variety of slopes. A similar conclusion has been drawn for bursts detected by BATSE (Shen et al. 2006), where the observed distribution of $p$, centered around $\langle p\rangle=2.50$ with $\sigma=0.49$, was found to be inconsistent with a universal value.
The second consideration regards the steepness of the $p$ slopes. In fact, the distribution of $p$ shows values generously extending at $p\sim$ 5 (up to 7 in one case). Such steep values of $p$ have been found also in other works in the literature. As an example, the synchrotron modeling of the prompt emission spectrum of GRB 180720B shown in Ronchi et al. (2020) returned a slope $p=4.8$ of the injected population of electrons (see their Fig. 5, SED I [0-35 s]). This in turn produced a soft photon spectrum ($\beta\sim-3.4$) at high energies, as confirmed also by the fit of the high-energy spectrum with a single power law. Such steep values of $p$ set strong constraints on the accelerations mechanism, either internal shocks (Bednarz & Ostrowski 1998; Achterberg et al. 2001; Lemoine & Pelletier 2003, see Sironi et al. 2015a for a recent review) or magnetic reconnection (Spruit et al. 2001; Drenkhahn & Spruit 2002), (Sironi & Spitkovsky 2014; Sironi et al. 2015b; Petropoulou et al. 2019). Indeed, the acceleration mechanism responsible for the injection of non-thermal electrons must be efficient in accelerating electrons to high energies, but it also should give rise to a steep electron energy distribution.
@@ 6.1 Comparison with previous detections of exponential cutoff
We here compare our results with earlier published works that found the presence of an exponential cutoff in the prompt emission spectrum. For this comparison, we exclude those works reporting a cutoff in the extra power–law components in the spectrum (e.g. see Ackermann et al. 2010, 2011, 2013; Abdo et al. 2009a, b), as we focus on the main prompt emission component, typically represented by the peaked Band-like spectral shape.
Fermi Large Area Telescope Team et al. 2012 analyzed 288 GRBs detected by GBM that do not show evidence of emission above 100 MeV. They compared the computed flux upper limits in the LAT energy band with the extrapolations of the spectral fits of GBM data, finding that 6 GRBs of the sample required the presence of a spectral cutoff between GBM and LAT data. However, lacking a direct evidence of a spectral cutoff, they set $E=100$ MeV as the cutoff energy for all 6 the bursts.
Tang et al. 2015 systematically searched for a high-energy spectral cutoff combining GBM and LAT data of 28 GRBs prompt emission spectra. They fitted the time-integrated prompt spectrum with the Band function with and without a high–energy cutoff. They found five GRBs showing a cutoff at energies between $\sim$13 and $\sim$62 MeV. Three of the five GRBs reported to have a cutoff are also present in our sample. Only in one of them, namely GRB 140206, we also found evidence of a spectral cutoff with $E_{\rm cutoff}=88.3^{+18.2}_{-12.5}$ MeV which is higher than the value $50.1\pm$6.8 MeV reported in (Tang et al. 2015). For the remaining 2 GRBs, we do not found any evidence of a cutoff and our best–fitting model is the 2SBPL. These discrepancies could be ascribed to the use of different fitting models (Band function in their analysis vs the set of SBPL/2SBPL in our work), although a more quantitative comparison should be performed to support this explanation.
@@ 6.2 Comparison with estimates of $\Gamma$ from pair-production opacity in the literature
In the work of Tang et al. 2015, the computation of pair-production opacity $\tau_{\gamma\gamma}$ is similar to the one in this work and also to the one presented in Lithwick & Sari (2001), namely assuming a simple one–zone model where the photon field in the emitting region is uniform, isotropic, and time independent. However, since the cutoff energies in the main spectral components they reported were below $\sim$ 60 MeV, in the majority of the cases their estimates of $\Gamma$ exceeded the maximum attainable values $\Gamma_{\rm max}$. This led the authors to use $\Gamma_{\rm max}$ as the actual bulk Lorentz factor for the majority of the bursts analyzed, finding values of $\Gamma$ between $\sim$ 50 and $\sim$150. These values are systematically lower than the values inferred in this work.
@@
For the remarkable cases of GRB 100724B and GRB 160509A, showing cutoff energies from $\sim$ few tens of MeV up to 150 MeV, Vianello et al. (2018) used two different time-dependent models to derive the values of the bulk Lorentz factors (i.e. an internal–shocks motivated model by Granot et al. 2008 and a photospheric model by Gill & Thompson 2014). The authors reported bulk Lorentz factors in the range $\Gamma\sim$ 100 – 400, which are consistent with our results.
@@
It is important to note that simplified one-zone models, as the one used in this work and in Lithwick & Sari (2001), Tang et al. (2015) and Fermi Large Area Telescope Team et al. (2012), may provide systematic differences in the inferred Lorentz factors as compared to more detailed time-dependent multi-zone models, where multiple emitting regions and the time-, space- and direction-dependent photon field are taken into account. Granot et al. (2008), and more recently Hascoët et al. (2012) and Gill & Granot (2018), have shown that such time-dependent models can yield inferred $\Gamma$ estimates that are reduced by a factor of $\sim$2 compared to estimates made using single-zone models. In the context of these time-dependent models, the estimates of $\Gamma$ presented in Fig. 5 should all be rescaled downward by a factor of $\sim$2. Moreover, Granot et al. (2008), investigating the opacity effects for impulsive emission between a given radius $R_{0}$ and $R_{0}+\Delta R$, found that the time-integrated spectrum should display a power-law tail (for impulsive emission with $\Delta R/R_{0}\lesssim 1$) instead of an exponential cutoff (for $\Delta R/R_{0}\gtrsim 1$ see also the more recent work of Dai et al. 2023) above a given cutoff energy. However, the signal-to-noise of spectral data in the MeV-GeV range typically offered by the current instruments makes it challenging to distinguish a break into a steep power-law from an exponential cutoff at high energies, although a direct comparison of these two models (broken power-law and exponential cutoff) on actual spectral data should be performed to investigate this. The model described in Granot et al. 2008, however, has been successfully applied to the cases of GRB 100724B and GRB 160509A in Vianello et al. (2018), where a spectral cutoff found using an empirical function (similarly to what done in this work) has been equally well fitted by it. For a discussion of single and multi-zone models, see also Zou et al. (2011).
@@
We note here that an alternative interpretation of the observed exponential cut-off is the maximum frequency for synchrotron emission, which reflects the maximum energy $\gamma_{\rm max}$ of the shock-accelerated particles distribution (Guilbert et al. 1983). The maximum energy $\gamma_{\rm max}$ is directly connected to the shock micro-physics conditions and it is typically estimated by equating the timescale for synchrotron cooling and the acceleration timescale. It is expected at much higher energies with respect to the cut-off energies found in this work, both for the electrons ($E^{\prime\rm max,syn}_{\rm e}\sim$ 50 MeV in the rest frame) and for the protons ($E^{\prime\rm max,syn}_{\rm p}\sim$ 93 GeV, see Ghisellini et al. 2020). We stress that this value is a limit and therefore, in principle, the exponential cut-off energies observed in this analysis could be explained as the maximum synchrotron frequency, provided that the distribution of accelerated particles spans a quite limited energy range from $\gamma_{\rm min}$ (given by $E_{\rm peak}$) to $\gamma_{\rm max}$ (corresponding to the cut-off energy $E_{\rm cutoff}$) and that the radiative losses are due to synchrotron emission. Under this interpretation, our results on the cut-off energies would provide only lower limits on the bulk Lorentz factor of the jet.
@@ 6.3 Comparison with $\Gamma$ derived from the afterglow onset time
The majority of the estimates of the bulk Lorentz factor $\Gamma$ for GRBs reported in the literature have been inferred from the measurement of the time of peak in the afterglow lightcurve (Molinari et al. 2007; Liang et al. 2010; Lü et al. 2012; Ghirlanda et al. 2012, 2018). In the standard fireball scenario, such a peak is commonly interpreted as due to the transition from the coasting (when the jet travels with a constant bulk Lorentz factor) to the deceleration phase (when the jet starts to decelerate due to the interaction with the interstellar medium), pinpointing the onset of the afterglow.
@@
Figure 7: Comparison of the distributions of bulk Lorentz factors of GRBs obtained from independent methods. The blue filled histogram represents the distribution of $\Gamma$ values derived in this work from the pair-production opacity argument applied to the prompt spectra of 9 GRBs showing a cutoff. The empty histograms show the distributions of $\Gamma$ values derived from the measurement of the afterglow onset time for a sample of 50 GRBs, as reported in Ghirlanda et al. 2018. The orange histogram represents the values of the bulk Lorentz factor derived assuming a constant medium density, while the green histogram refers to the wind–like medium case. These distributions are built taking into account measurements only (not lower limits), for both methods.
@@
Overall, the values of $\Gamma$ derived in this work are consistent with those derived from the afterglow onset reported in Ghirlanda et al. (2018). Our results are well included in the 68% of the distribution of $\Gamma$ derived from the afterglow onset time assuming an homogeneous medium. A K–S test performed between the constant ISM case and pair–production $\Gamma$ distributions returns a $p$-value = 0.64, indicating that they are likely drawn from the same underlying distribution of bulk Lorentz factors. Our results are instead located in the higher tail of the distribution related to the wind–like medium, and a K–S test returns a $p$-value = 0.003. Since in the wind case the assumed density of the circumburst medium is higher, the $\Gamma$ values related to the wind–like case are typically smaller than the ones related to the homogeneous case. Keeping fixed all the other parameters in Eq. 7, smaller values of $\Gamma$ would imply smaller values of the cutoff energy $E_{\rm cutoff}$, which in turn implies that the high energy prompt emission may not be detected in the LAT energy band. In turn, this may indicate a selection effect in favor of faster GRBs when analyzing GRBs detected by Fermi/LAT. On the other hand, there may be a bias also on the right side of the distribution, namely towards higher values of the bulk Lorentz factor. The distribution of $\Gamma$ may extend towards higher values, as indeed already suggested by the distribution of lower-limits on $\Gamma$ (red dashed histogram in Fig. 5), which extends up to $\Gamma\sim$750. However, keeping fixed all the other parameters of Eq. 7, such higher values of $\Gamma$ imply higher values of $E_{\rm cutoff}$, which cannot be observed with the current instruments.
@@
Since the $\Gamma$ values derived from the afterglow onset time depends on the medium density, this comparison may be affected by a different choice for this parameter (and for the wind velocity and mass loss rate). For example, in order to reproduce the $\Gamma$ distribution found in our analysis (blue filled histogram), the distribution derived from the afterglow onset for the wind case (green empty histogram) should be moved towards higher values of $\Gamma$ and this implies to assume smaller densities of the external medium. This is more relevant for the wind case ($\Gamma\propto n^{-1/4}$) than in the case of homogeneous ISM ($\Gamma\propto n^{-1/8}$).

### v2

@@ 5.2 Estimate of p
Figure 6: Distribution of the slope $p$ of the nonthermal population of accelerated particles for the sample of 22 GRBs analyzed in this work. This distribution has been derived from the $\beta$ distribution shown in Fig. 4, by using the relation $p=-2-2\beta$ or, equivalently, $p=2\beta_{e}$ (i.e., assuming synchrotron emission in the fast cooling regime). The median value of $p$ along with its error bar on $p$ values are represented with the black dot and solid line, respectively. For comparison, the distribution of $p$ derived for the Fermi/GBM catalog bursts under the same assumption is also plotted (red dashed histogram).
There are two main considerations that we can draw from these results. The first one regards the width of the $p$ distribution. Fig. 6 shows that the distribution of $p$ is quite broad, with a $\sigma=1.12$. The majority of $p$ values are clustered between 2 and 3.5, but there is also a non-negligible tail extending toward steep values. This result may suggest that the acceleration mechanism does not produce relativistic particles distributed as a power–law with a unique slope, but rather with a quite wide variety of slopes. A similar conclusion has been drawn for bursts detected by BATSE (Shen et al. 2006), where the observed distribution of $p$, centered around $\langle p\rangle=2.50$ with $\sigma=0.49$, was found to be inconsistent with a universal value.
The second consideration regards the steepness of the $p$ slopes. In fact, the distribution of $p$ shows values generously extending at $p\sim$ 5 (up to 7 in one case). Such steep values of $p$ have been found also in other works in the literature. As an example, the synchrotron modeling of the prompt emission spectrum of GRB 180720B shown in Ronchi et al. (2020) returned a slope $p=4.8$ of the injected population of electrons (see their Fig. 5, SED I [0-35 s]). This in turn produced a soft photon spectrum ($\beta\sim-3.4$) at high energies, as confirmed also by the fit of the high-energy spectrum with a single power law. Such steep values of $p$ set strong constraints on the accelerations mechanism, either internal shocks (Bednarz & Ostrowski 1998; Achterberg et al. 2001; Lemoine & Pelletier 2003, see Sironi et al. 2015a for a recent review) or magnetic reconnection (Spruit et al. 2001; Drenkhahn & Spruit 2002), (Sironi & Spitkovsky 2014; Sironi et al. 2015b; Petropoulou et al. 2019). Indeed, the acceleration mechanism responsible for the injection of nonthermal electrons must be efficient in accelerating electrons to high energies, but it also should give rise to a steep electron energy distribution.
@@ 6.1 Comparison with previous detections of exponential cutoff
We here compare our results with earlier published works that found the presence of an exponential cutoff in the prompt emission spectrum. For this comparison, we exclude those works reporting a cutoff in the extra power–law components in the spectrum (e.g., see Ackermann et al. 2010, 2011, 2013; Abdo et al. 2009a, b), as we focus on the main prompt emission component, typically represented by the peaked Band-like spectral shape.
Fermi Large Area Telescope Team et al. 2012 analyzed 288 GRBs detected by GBM that do not show evidence of emission above 100 MeV. They compared the computed flux upper limits in the LAT energy band with the extrapolations of the spectral fits of GBM data, finding that 6 GRBs of the sample required the presence of a spectral cutoff between GBM and LAT data. However, lacking a direct evidence of a spectral cutoff, they set $E=100$ MeV as the cutoff energy for all 6 bursts.
Tang et al. 2015 systematically searched for a high-energy spectral cutoff combining GBM and LAT data of 28 GRBs prompt emission spectra. They fit the time-integrated prompt spectrum with the Band function with and without a high–energy cutoff. They found five GRBs showing a cutoff at energies between $\sim$13 and $\sim$62 MeV. Three of the five GRBs reported to have a cutoff are also present in our sample. Only in one of them, namely GRB 140206, we also found evidence of a spectral cutoff with $E_{\rm cutoff}=88.3^{+18.2}_{-12.5}$ MeV which is higher than the value $50.1\pm$6.8 MeV reported in (Tang et al. 2015). For the remaining 2 GRBs, we do not find any evidence of a cutoff and our best–fitting model is the 2SBPL. These discrepancies could be ascribed to the use of different fitting models (Band function in their analysis vs the set of SBPL/2SBPL in our work), although a more quantitative comparison should be performed to support this explanation.
@@ 6.2 Comparison with estimates of $\Gamma$ from pair-production opacity in the literature
In the work of Tang et al. 2015, the computation of pair-production opacity $\tau_{\gamma\gamma}$ is similar to the one in this work and also to the one presented in Lithwick & Sari (2001), namely assuming a simple one–zone model where the photon field in the emitting region is uniform, isotropic, and time-independent. However, since the cutoff energies in the main spectral components they reported were below $\sim$ 60 MeV, in the majority of the cases their estimates of $\Gamma$ exceeded the maximum attainable values $\Gamma_{\rm max}$. This led the authors to use $\Gamma_{\rm max}$ as the actual bulk Lorentz factor for the majority of the bursts analyzed, finding values of $\Gamma$ between $\sim$ 50 and $\sim$150. These values are systematically lower than the values inferred in this work.
@@
For the remarkable cases of GRB 100724B and GRB 160509A, showing cutoff energies from $\sim$ few tens of MeV up to 150 MeV, Vianello et al. (2018) used two different time-dependent models to derive the values of the bulk Lorentz factors (i.e., an internal–shocks motivated model by Granot et al. 2008 and a photospheric model by Gill & Thompson 2014). The authors reported bulk Lorentz factors in the range $\Gamma\sim$ 100 – 400, which are consistent with our results.
@@
It is important to note that simplified one-zone models, as the one used in this work and in Lithwick & Sari (2001), Tang et al. (2015) and Fermi Large Area Telescope Team et al. (2012), may provide systematic differences in the inferred Lorentz factors as compared to more detailed time-dependent multi-zone models, where multiple emitting regions and the time-, space- and direction-dependent photon field are taken into account. Granot et al. (2008), and more recently Hascoët et al. (2012) and Gill & Granot (2018), have shown that such time-dependent models can yield inferred $\Gamma$ estimates that are reduced by a factor of $\sim$2 compared to estimates made using single-zone models. In the context of these time-dependent models, the estimates of $\Gamma$ presented in Fig. 5 should all be rescaled downward by a factor of $\sim$2. Moreover, Granot et al. (2008), investigating the opacity effects for impulsive emission between a given radius $R_{0}$ and $R_{0}+\Delta R$, found that the time-integrated spectrum should display a power-law tail (for impulsive emission with $\Delta R/R_{0}\lesssim 1$) instead of an exponential cutoff (for $\Delta R/R_{0}\gtrsim 1$ see also the more recent work of Dai et al. 2023) above a given cutoff energy. However, the signal-to-noise of spectral data in the MeV-GeV range typically offered by the current instruments makes it challenging to distinguish a break into a steep power-law from an exponential cutoff at high energies, although a direct comparison of these two models (broken power-law and exponential cutoff) on actual spectral data should be performed to investigate this. In the work of Vianello et al. (2018), where a spectral cutoff has been found in GRB 100724B and GRB 160509A using an empirical function (similar to what is done in this work) and also the model described in Granot et al. 2008, the authors tested whether a power-law after the cutoff energy provide a better fit than the exponential, and concluded that, despite the increased complexity, this is never the case, for both GRBs, suggesting that the shape of the spectrum after the cutoff appears to be curved. For a discussion of single and multi-zone models, see also Zou et al. (2011).
@@
We note here that an alternative interpretation of the observed exponential cutoff would be the maximum frequency for synchrotron emission, which reflects the maximum energy $\gamma_{\rm max}$ of the shock-accelerated particles distribution (Guilbert et al. 1983). The maximum energy $\gamma_{\rm max}$ is directly connected to the shock micro-physics conditions and it is typically estimated by equating the timescale for synchrotron cooling and the acceleration timescale. It is expected at much higher energies with respect to the cutoff energies found in this work, both for the electrons ($E^{\prime\rm max,syn}_{\rm e}\sim$ 50 MeV in the rest frame) and for the protons ($E^{\prime\rm max,syn}_{\rm p}\sim$ 93 GeV, see Ghisellini et al. 2020). We stress that this value is a limit and therefore, in principle, the exponential cut-off energies observed in this analysis could be explained as the maximum synchrotron frequency, provided that the distribution of accelerated particles spans a quite limited energy range from $\gamma_{\rm min}$ (given by $E_{\rm peak}$) to $\gamma_{\rm max}$ (corresponding to the cutoff energy $E_{\rm cutoff}$) and that the radiative losses are due to synchrotron emission. Under this interpretation, our results on the cutoff energies would provide only lower limits on the bulk Lorentz factor of the jet.
@@ 6.3 Comparison with $\Gamma$ derived from the afterglow onset time
The majority of the estimates of the bulk Lorentz factor $\Gamma$ for GRBs reported in the literature have been inferred from the measurement of the time of the peak in the afterglow light curve (Molinari et al. 2007; Liang et al. 2010; Lü et al. 2012; Ghirlanda et al. 2012, 2018). In the standard fireball scenario, such a peak is commonly interpreted as due to the transition from the coasting (when the jet travels with a constant bulk Lorentz factor) to the deceleration phase (when the jet starts to decelerate due to the interaction with the interstellar medium), pinpointing the onset of the afterglow.
@@
Fig. 7 shows the distributions of the values of the bulk Lorentz factors inferred from the two methods (taking into account only the measurements and not the lower–limits). The filled blue histogram represents the estimates of $\Gamma$ derived from the $\gamma\gamma$ opacity argument applied in this work, while the empty histograms represent the values derived from the measurement of the afterglow onset time as reported in Ghirlanda et al. 2018, with the orange (green) color representing the homogeneous (wind-like) medium. For the homogeneous medium case, the authors assumed a typical medium density of $n_{\rm 0}=1$ cm-3. For the wind-like case, they used a normalization of $n_{\rm 0}=10^{35}\dot{M}_{-5}\,v_{\rm w,3}^{-1}$ cm-3, following the relation $n(r)=n_{\rm 0}r^{-2}=\dot{M}/(4\pi m_{\rm p}v_{\rm w}r^{2})$, where typical values of the mass-loss rate $\dot{M}=10^{-5}M_{\odot}$ yr-1 and of the wind velocity $v_{\rm w}=10^{3}$ km s-1 have been assumed (Chevalier & Li 2000). We stress, however, that we are comparing the distributions assuming that all GRBs in Ghirlanda et al. 2018 have an ISM medium (orange distribution) or a wind (green distribution) medium, while of course for each burst only one medium type, and by extension only one value of $\Gamma$, is relevant, and a detailed analysis of the individual broad-band afterglow emission should be performed to distinguish the adequate medium for each burst.
Figure 7: Comparison of the distributions of bulk Lorentz factors of GRBs obtained from independent methods. The blue-filled histogram represents the distribution of $\Gamma$ values derived in this work from the pair-production opacity argument applied to the prompt spectra of 9 GRBs showing a cutoff. The empty histograms show the distributions of $\Gamma$ values derived from the measurement of the afterglow onset time for a sample of 50 GRBs, as reported in Ghirlanda et al. 2018. The orange histogram represents the values of the bulk Lorentz factor derived assuming a constant medium density, while the green histogram refers to the wind–like medium case. These distributions are built taking into account measurements only (not lower limits), for both methods.
@@
Overall, the values of $\Gamma$ derived in this work are consistent with those derived from the afterglow onset reported in Ghirlanda et al. (2018). Our results are well included in the 68% of the distribution of $\Gamma$ derived from the afterglow onset time assuming a homogeneous medium. A K–S test performed between the constant ISM case and pair–production $\Gamma$ distributions returns a $p$-value = 0.64, indicating that they are likely drawn from the same underlying distribution of bulk Lorentz factors. Our results are instead located in the higher tail of the distribution related to the wind–like medium, and a K–S test returns a $p$-value = 0.003. Since in the wind case the assumed density of the circumburst medium is higher, the $\Gamma$ values related to the wind–like case are typically smaller than the ones related to the homogeneous case. Keeping fixed all the other parameters in Eq. 7, smaller values of $\Gamma$ would imply smaller values of the cutoff energy $E_{\rm cutoff}$, which in turn implies that the high energy prompt emission may not be detected in the LAT energy band. In turn, this may indicate a selection effect in favor of faster GRBs when analyzing GRBs detected by Fermi/LAT. On the other hand, there may be a bias also on the right side of the distribution, namely toward higher values of the bulk Lorentz factor. The distribution of $\Gamma$ may extend toward higher values, as indeed already suggested by the distribution of lower-limits on $\Gamma$ (red dashed histogram in Fig. 5), which extends up to $\Gamma\sim$750. However, keeping fixed all the other parameters of Eq. 7, such higher values of $\Gamma$ imply higher values of $E_{\rm cutoff}$, which cannot be observed with the current instruments.
@@
Since the $\Gamma$ values derived from the afterglow onset time depends on the medium density, this comparison may be affected by a different choice for this parameter (and for the wind velocity and mass loss rate). For example, in order to reproduce the $\Gamma$ distribution found in our analysis (blue filled histogram), the distribution derived from the afterglow onset for the wind case (green empty histogram) should be moved toward higher values of $\Gamma$ and this implies to assume smaller densities of the external medium. This is more relevant for the wind case ($\Gamma\propto n^{-1/4}$) than in the case of homogeneous ISM ($\Gamma\propto n^{-1/8}$).